In the game of roulette, there are 38 total numbers on a Roulette wheel. They are numbered 1-36 with the last two being 0 and 00. One of the bets you can make is called dozens, where you bet on a particular group of 12 numbers (either 1-12, 13-24, or 25-36) occurring. You bet $100 on one of these dozens to win $300. All 38 numbers on the wheel are equally likely. What is the probability that you win this bet?
Answer #1

as All 38 numbers on the wheel are equally likely and you are betting on 12 of them;

probability that you win this bet =12/38=6/19